# Earth Observation Foundation Model PhilEO: Pretraining on the MajorTOM and FastTOM Datasets

### v3

@@ Abstract
Today, Earth Observation (EO) satellites generate massive volumes of data. To fully exploit this, it is essential to pretrain Foundation Models (FMs) on large unlabeled datasets, enabling efficient fine-tuning for downstream tasks with minimal labeled data. We study scaling-up FMs, where scaling refers to incorporating a larger pretraining dataset that includes both oceans and ice in addition to land (MajorTOM, 23TB), a larger model in terms of number of parameters (200M), and an architecture that achieves top performance, efficiency, and low computational complexity, being capable of handling images covering large areas. We develop various models using different architectures, including U-Net Convolutional Neural Networks (CNN), Vision Transformers (ViT), and Mamba State-Space Models (SSM), and different numbers of parameters. We evaluate the FLoating-point OPerations (FLOPs) needed by the models. ViT, due to attention, suffers from quadratic complexity, while Mamba achieves linear complexity, scaling-up better. We present the scaling of the model PhilEO on MajorTOM, covering the vast majority of Earth, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. We fine-tune the models on the PhilEO Bench for different downstream tasks: roads, buildings, and land cover. For all n-shots for roads, the Geo-Aware U-Net 44M-23T model outperforms U-Net 44M-0.5T. For most n-shots for roads and buildings, U-Net 200M-2T outperforms the other models. We show that using Mamba models, we achieve comparable results on the downstream tasks, with less computational expenses. We also compare with the recent FM TerraMind which we evaluate on the PhilEO Bench.
Index Terms:
Earth Observation, Foundation Models, remote sensing, Vision Transformer (ViT), Mamba state-space model.
@@ I Introduction
In this work, our main aim is to study the scaling-up of EO FMs, where the scaling-up refers to incorporating a larger pretraining dataset that includes both oceans and ice in addition to land, for example the MajorTOM dataset 23TB (Sentinel-2 L2A, Core-S2L2A), a larger model in terms of number of parameters, for example 200M, and an architecture that is suitable for scaling-up in terms of performance, efficiency, speed, and computational complexity, being capable of handling high-resolution data and images that cover large areas. Regarding the latter, i.e. regarding the architecture we select to scale-up, we note that the Vision Transformer (ViT), because of attention, suffers from quadratic complexity with respect to the image token sequence length. On the contrary, the recent Mamba S6 State-Space Model (SSM) achieves linear complexity and scales more favorably to high-resolution data and images that cover large geographical areas [30]. In this way, global correlations and long-range dependencies within the EO image that covers a large area can be effectively and efficiently modelled. Moreover, U-Net Convolutional Neural Network (CNN) models also have linear complexity, but can not model global correlations and long-range dependencies. In addition, in this work, scaling-up also refers to including the time aspect in the pretraining, for example seasonality and the four seasons of summer, winter, autumn, and fall, as well as additional modalities, for example Sentinel-1 (S-1) SAR data in addition to Sentinel-2 (S-2) multi-spectral images.
Our scaling-up study. In this paper, we present the scaling-up of our recently proposed model PhilEO on the unlabeled 23TB dataset MajorTOM [15], which covers the vast majority of the Earth’s surface, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. FastTOM is a task-specific dataset for land only. In the literature, there exist FMs only for oceans or ice, for example HydroFM [25], but we are interested in Earth Observation and not only in Land Observation or Ocean Observation (or Ice Observation). To this end, in this work, we develop and study various models with different numbers of parameters, and also with different architectures. We evaluate and examine the FLoating-point OPerations (FLOPs) that the models need [65]. We fine-tune the models on the PhilEO Bench [3] for road density estimation, building density pixel-wise regression, and land cover semantic segmentation, and we evaluate their performance.
@@
  • We train our model on the pretraining dataset MajorTOM 23TB which includes all regions, and the performance is competitive versus models pretrained on more specialized datasets (i.e. substantially smaller datasets that only include land). The additional data of oceans and ice on average do not decrease the performance on land-focused downstream tasks. These results indicate that large GFMs trained on global datasets for a wider variety of downstream tasks can be useful for downstream applications that only require a subset of the information included in their training.
@@
We scale the pretraining from 0.5TB to 23TB, as well as to 2TB, where the 23TB are general-purpose EO data. We use different number of model parameters and architectures [^1] We release the code of our model in: http://github.com/ESA-PhiLab/PhilEO-MajorTOM.. We show that for road density regression, for the $n$-shot $n=500$, larger pretraining data reduces the final downstream task’s Root Mean Squared Error (RMSE), resulting in a percentage improvement $12.0\%$. In addition to pretraining dataset scaling, we also study the scaling-up of the model parameters and the architecture, comparing CNN- and ViT-based models, as well as Mamba. We evaluate the effectiveness of the scaling-up on the PhilEO Bench, covering both road and building density estimation, as well as land cover mapping.
@@
Organization of this paper. The rest of the paper is organized as follows. Section II provides background on the PhilEO Geospatial Foundation Model (GFM) and the PhilEO Bench evaluation, as well as the datasets MajorTOM 23TB and FastTOM 2TB. Section III describes the proposed model. Section IV presents the related work. We evaluate the models on the PhilEO Bench evaluation framework in Sec. V for road and building density estimation, and land cover mapping. We present the results of the PhilEO MajorTOM 23TB and FastTOM 2TB models, as well as the results of the PhilEO Geo-Aware U-Net, the PhilEO ViT UPerNet 100M MajorTOM and FastTOM, and the PhilEO 200M FastTOM.
@@ II Background
PhilEO GFM. The PhilEO model [3] employs a combination of masked reconstruction and geo-location estimation as pretext tasks to train on the PhilEO Globe 0.5TB S2L2A dataset, which contains only land, excluding oceans and ice. PhilEO Globe also includes four time steps: $4$ seasons from the same year. The architecture is a modified U-Net CNN model.
@@ III-A PhilEO MajorTOM 23TB and PhilEO FastTOM 2TB
In this subsection, we note that our main contribution is pretraining on the FastTOM and MajorTOM datasets. For the pretraining, the models we develop, implement, and train have three output heads [3], one for reconstruction with the Mean Squared Error (MSE) loss where we mask a percentage of the input image during training, one for estimating the geo-location longitude and latitude where cosine and sine values are used, and one for predicting the Köppen-Geiger climate zone classes [3], [67]. “Geo-Aware” (for example, in the PhilEO Geo-Aware U-Net model) refers to these three output heads, i.e. to the reconstruction objective, the geo-location estimation, and the Köppen-Geiger climate zone prediction.
@@ III-C Mamba 2D Models
For the Mamba 2D models which we will also examine in Sec. V, they are based on continuous-time dynamics [30], [33]. We denote the state by $h$, and for the linear dynamics (i.e. similar to the Kalman filter), we have:
@@
where $\textbf{W}_{\Delta}\in\mathbb{R}^{1\times D}$ is the learnable linear projection (i.e., fully-connected (FC) layer) for making the time step $\Delta$ input-dependent [30], [63]. In addition, $\textbf{S}_{\Delta}$ is a function of the output of the learnable linear projection, $\textbf{x}\textbf{W}_{\Delta}^{T}$. Moreover, in (7) and (8), $\tau_{\Delta}$ is the softplus function. The $\text{Broadcast}_{D}(.)$ operation broadcasts the result to all the $D$ dimensions, i.e. from $1$ to $D$, as $\textbf{S}_{\Delta}\in\mathbb{R}^{B\times L\times D}$.
@@
ViT models have significantly less inductive bias compared to CNN U-Net models. This is because ViT attention correlates everything with everything else. Moreover, in self-attention, image token order does not matter [9], [30]. In addition, Mamba models have more image-inductive bias compared to ViT models, as in Mamba S6 SSM models, the image token order is important and matters. In addition, we also note that our main contribution in this subsection is including Mamba SSM 2D models in the PhilEO Bench (i.e. see Sec. V-D).
@@ III-D The S12-PhilEOBench evaluation framework
In this work, we also extend the PhilEO Bench dataset [3] to include co-located S-1 SAR data in addition to S-2 multi-spectral images. In Figs. 2 and 3, we show example images [^3] The S-1 SAR PhilEO Bench data can be found in http://huggingface.co/datasets/NikolaosDionelis2023/s1-phileobench/tree/main. In addition, we note that the S-2 multi-spectral PhilEO Bench data can be found in http://huggingface.co/datasets/NikolaosDionelis2023/s2-phileobench/tree/main. from the global dataset S12-PhilEOBench (for example, Fig. 2 is from East Africa). For the same S-2 and S-1 images, we have three types of labels: road density pixel-wise regression (continuous real-valued output), building density regression, and land cover pixel-wise classification (with $11$ classes).
@@ V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
For the evaluation of the models, continuing the discussion in the preceding paragraph, we count the FLOPs that are needed by the different models [65]. The lower the FLOPs, the faster the model, i.e. the better (less operations). We measure the FLOPs in Python using the DeepSpeed library. Here, we also note that FLOPs are different from Floating-point operations Per Second (FPS). To be machine GPU agnostic, we use FLOPs and we do not measure time (i.e. seconds). In our FLOPs profiling analysis, we also include the models TerraMind [29] and Prithvi-EO-2.0 [23] (see Sec. V-E).
@@
Sentinel-2, L2A data. For the models we develop, implement, train, and evaluate that are based on the EO FM PhilEO [3], we use images of size 128128x128128, and also $10$ spectral bands (channels), i.e. multi-spectral satellite Sentinel-2 data. More specifically, as in [3], we use the $10$ and $20$ metre spatial resolution spectral bands from the satellite Sentinel-2777http://sentiwiki.copernicus.eu/web/s2-mission#S2Mission-SpatialResolutionS2-Mission-Spatial-Resolutiontrue.
@@ V-C Comparison of scaling-up for U-Net and ViT models
Jointly increasing the pretraining dataset size and the model size. In this paper, we focus not only on scaling-up the volume of the pretraining data (i.e. to 23TB) [15], but also on scaling the number of parameters and the architecture, also studying the transition from CNN-based U-Net models to ViT backbones combined with the UPerNet decoder, as well as to Mamba for images. Here, the recent Mamba S6 SSM is a potential effective paradigm-shift solution [30], as it combines global dependencies modeling and linear computational scaling, and this is important. The self-attention mechanism of the ViT is not sensitive to the token order, while on the contrary, the selectivity mechanism of Mamba (i.e. the selective scan) is sensitive to the token order. Notably, it needs a large hidden state dimension as it has an exponentially decaying recall accuracy based on the distance from the current token [36]. In addition, ViT has quadratic complexity with respect to the image token sequence length, due to the attention Query, Key, Value mechanism, which is based on the scaled dot product similarity measure, while the 2D Mamba S6 SSM model is faster and more efficient as it achieves linear complexity with respect to the image token sequence length [30], [31].
@@
Figure 18: Road density pixel-wise regression in RMSE for the models ViT and Mamba, and comparison to the TerraMind and Prithvi-EO-2.0 models. In the plot, the green line (Prithvi-EO-2.0) is below the pink line (TerraMind).
@@ VI Conclusion
In this paper, we also compared with Mamba models (see Sec. III-C, as well as Sec. V-D), and this also one of our main contributions. Furthermore, we think our FLOPs profiling analysis, comparison, and benchmarking (i.e. in Sec. V, as well as in Figs. 8 and 9) opens the road to fairly evaluating different models (including GFMs) and will inspire others to adopt and use this real-world setting. Finally, it was a major achievement to process the 23TB of MajorTom on 32 cores of the Leonardo supercomputer davinci-1, in many days. This would have taken many more days, and even more months, on a single GPU. As future work, we plan to use JEPA, as well as add multi-temporal capability (i.e. multiple time steps) [58], and multi-modality. Regarding the latter, we will add S-1 in addition to S-2 in the pretraining and not only in the S12-PhilEOBench (i.e. see Sec. III-D). As future work, we will also use Mamba 2D models with larger images, for example 512512x512512 (or even 10241024x10241024), as well as perform knowledge distillation of models to the same number of FLOPs (e.g., for On-Board AI) and compare and benchmark. Finally, as future work, we plan to also examine, compare with, and try to include in the model pipeline the recent AlphaEarth Foundations Satellite Embeddings [^8] http://deepmind.google/discover/blog/alphaearth-foundations-helps-map-our-planet-in-unprecedented-detail/. These are geo-located embeddings at 1010m resolution, and they also need the longitude and latitude information to be used in the downstream tasks.

### v4

@@ Abstract
Today, Earth Observation (EO) satellites generate massive volumes of data. To fully exploit this, it is essential to pretrain Foundation Models (FMs) on large unlabeled datasets, enabling efficient fine-tuning for downstream tasks with minimal labeled data. In this paper, we study scaling-up FMs: we train our models on the pretraining dataset MajorTOM 23TB which includes all regions, and the performance on average is competitive versus models pretrained on more specialized datasets which are substantially smaller and include only land. The additional data of oceans and ice do not decrease the performance on land-focused downstream tasks. These results indicate that large FMs trained on global datasets for a wider variety of downstream tasks can be useful for downstream applications that only require a subset of the information included in their training. The second contribution is the exploration of U-Net Convolutional Neural Network (CNN), Vision Transformers (ViT), and Mamba State-Space Models (SSM) as FMs. U-Net captures local correlations amongst pixels, while ViT and Mamba capture local and distant correlations. We develop various models using different architectures, including U-Net, ViT, and Mamba, and different number of parameters. We evaluate the FLoating-point OPerations (FLOPs) needed by the models. We fine-tune on the PhilEO Bench for different downstream tasks: roads, buildings, and land cover. For most n-shots for roads and buildings, U-Net 200M-2T outperforms the other models. Using Mamba, we achieve comparable results on the downstream tasks, with less computational expenses. We also compare with the recent FM TerraMind which we evaluate on PhilEO Bench.
@@ I Introduction
In this work, our main aim is to study the scaling-up of EO FMs. Here, scaling-up refers to: i) incorporating a larger pretraining dataset that includes both oceans and ice in addition to land, i.e. the MajorTOM dataset 23TB (Sentinel-2 L2A, Core-S2L2A) [15], ii) a larger model in terms of number of parameters, i.e. 200M, and iii) an architecture that is suitable for scaling-up in terms of performance, efficiency, speed, and computational complexity, being capable of handling high-resolution data and images that cover large areas. Regarding the latter, i.e. regarding the architecture we select to scale-up, we note that the Vision Transformer (ViT), because of attention, suffers from quadratic complexity with respect to the image token sequence length. On the contrary, the recent Mamba S6 State-Space Model (SSM) achieves linear complexity and scales more favorably to high-resolution data and images that cover large geographical areas [30]. In this way, global correlations and long-range dependencies within the EO image that covers a large area can be effectively and efficiently modelled. Moreover, U-Net Convolutional Neural Network (CNN) models also have linear complexity, but can not model global correlations and long-range dependencies.
We use and train our recently proposed model PhilEO [3]. MajorTOM covers the vast majority of the Earth’s surface. In this paper, we also create and use the specialized subset FastTOM 2TB that does not include oceans and ice. FastTOM is a task-specific dataset for land only. In the literature, there exist FMs only for oceans or ice, for example HydroFM [25], but we are interested in Earth Observation and not only in Land Observation or Ocean Observation (or Ice Observation). To this end, in this work, we develop and study various models with different numbers of parameters, and also with different architectures. We evaluate and examine the FLoating-point OPerations (FLOPs) that the models need [65]. We fine-tune the models on the PhilEO Bench [3] for road density estimation, building density pixel-wise regression, and land cover semantic segmentation, and we evaluate their performance.
@@
  • We train our models on the pretraining dataset MajorTOM 23TB which includes all regions, and the performance is competitive versus models pretrained on more specialized datasets (i.e. substantially smaller datasets that only include land). The additional data of oceans and ice on average do not decrease the performance on land-focused downstream tasks. These results indicate that large GFMs trained on global datasets for a wider variety of downstream tasks can be useful for downstream applications that only require a subset of the information included in their training.
@@
Organization of this paper. The rest of the paper is organized as follows. Section II provides background on the PhilEO Geospatial Foundation Model (GFM) and the PhilEO Bench evaluation, as well as the datasets MajorTOM 23TB and FastTOM 2TB. Section III describes the proposed model [^1] We release the code of our model in: http://github.com/ESA-PhiLab/PhilEO-MajorTOM.. Section IV presents the related work. We evaluate the models on the PhilEO Bench evaluation framework in Sec. V for road and building density estimation, and land cover mapping. We present the results of the PhilEO MajorTOM 23TB and FastTOM 2TB models, as well as the results of the PhilEO Geo-Aware U-Net, the PhilEO ViT UPerNet 100M MajorTOM and FastTOM, and the PhilEO 200M FastTOM.
@@ II Background
PhilEO GFM. The PhilEO model [3] employs a combination of masked reconstruction and geo-location estimation as pretext tasks to train on the PhilEO Globe 0.5TB S2L2A dataset, which contains only land, excluding oceans and ice. The architecture is a modified U-Net CNN model. PhilEO Globe also includes four time steps: $4$ seasons from the same year. Regarding the four time steps in PhilEO Globe 0.5TB, we note that in this work, scaling-up also refers to including the time aspect in the pretraining, for example seasonality and the four seasons of summer, winter, autumn, and fall.
@@ III-A PhilEO MajorTOM 23TB and PhilEO FastTOM 2TB
For the pretraining, the models we develop and train have three output heads [3], one for reconstruction with the Mean Squared Error (MSE) loss where we mask a percentage of the input image during training, one for estimating the geo-location longitude and latitude where cosine and sine values are used, and one for predicting the Köppen-Geiger climate zone classes [3], [67]. “Geo-Aware” (for example, in the PhilEO Geo-Aware U-Net model) refers to these three output heads, i.e. to the reconstruction objective, the geo-location estimation [^3] http://huggingface.co/datasets/NikolaosDionelis2023/data_statics/tree/main, and the Köppen-Geiger climate zone prediction.
In this subsection, we note that our main contribution is pretraining on the MajorTOM and FastTOM [^4] http://huggingface.co/datasets/NikolaosDionelis2023/Subset_FastTOM/tree/main datasets. We scale the pretraining from 0.5TB to 23TB, as well as to 2TB, where the 23TB are general-purpose EO data. We use different number of model parameters and architectures. We show that for road density regression, for the $n$-shot $n=500$, larger pretraining data reduces the final downstream task’s Root Mean Squared Error (RMSE), resulting in a percentage improvement $12.0\%$. In addition to pretraining dataset scaling, we also study the scaling-up of the model parameters and the architecture, comparing CNN- and ViT-based models, as well as Mamba. We evaluate the effectiveness of the scaling-up on the PhilEO Bench, covering both road and building density estimation, as well as land cover mapping.
@@ III-C Mamba 2D Models
For the Mamba 2D models, which we will also examine in Sec. V, they are based on continuous-time dynamics [30], [33]. We denote the state by $h$, and for the linear dynamics (i.e. similar to the Kalman filter), we have:
@@
where $\textbf{W}_{\Delta}\in\mathbb{R}^{1\times D}$ is the learnable linear projection (i.e., fully-connected (FC) layer) for making the time step $\Delta$ input-dependent [30], [63]. In addition, $\textbf{S}_{\Delta}$ is a function of the output of the learnable linear projection, $\textbf{x}\textbf{W}_{\Delta}^{T}$. Moreover, in (7) and (8), $\tau_{\Delta}$ is the softplus function. The $\text{Broadcast}_{D}(.)$ operation broadcasts the result to all the $D$ dimensions, i.e. from $1$ to $D$, as $\textbf{S}_{\Delta}\in\mathbb{R}^{B\times L\times D}$. Using the parameter $\Delta$ as well as the equations (1)-(8), Mamba SSM models also use the Zero-Order Hold (ZOH) condition so as to treat each discrete input sample as a constant value that is held for a specific time interval until the next data sample is introduced [30], [69].
@@
ViT models have significantly less inductive bias compared to CNN U-Net models. This is because ViT attention correlates everything with everything else. Moreover, in self-attention, image token order does not matter [9], [30]. In addition, Mamba models have more image-inductive bias compared to ViT models, as in Mamba S6 SSM models, the image token order is important and matters. In addition, we also note that our main contribution in this subsection is including Mamba SSM 2D models in the PhilEO Bench (i.e. see Sec. V-D), examining their performance as they can capture global correlations and have linear (and not quadratic) complexity.
@@ III-D The S12-PhilEOBench evaluation framework
In addition, in this paper, scaling-up also refers to including additional modalities, for example Sentinel-1 (S-1) SAR data in addition to Sentinel-2 (S-2) multi-spectral images. To this end, in this work, we also extend the PhilEO Bench dataset [3] to include co-located S-1 SAR data in addition to S-2 multi-spectral images. In Figs. 2 and 3, we show example images [^5] The S-1 SAR PhilEO Bench data can be found in http://huggingface.co/datasets/NikolaosDionelis2023/s1-phileobench/tree/main. In addition, we note that the S-2 multi-spectral PhilEO Bench data can be found in http://huggingface.co/datasets/NikolaosDionelis2023/s2-phileobench/tree/main. from the global dataset S12-PhilEOBench (for example, Fig. 2 is from East Africa). For the same S-2 and S-1 images, we have three types of labels: road density pixel-wise regression (continuous real-valued output), building density regression, and land cover pixel-wise classification (with $11$ classes).
@@ V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
For the evaluation of the models, continuing the discussion in the preceding paragraph, we count the FLOPs needed by the different models [65]. The lower the FLOPs, the faster the model, i.e. the better (less operations). We measure the FLOPs in Python using the DeepSpeed library. We note that similar values are also obtained using the Python libraries fvcore and calflops. We also note that FLOPs are different from Floating-point operations Per Second (FPS). To be machine GPU agnostic, we use FLOPs and we do not measure time (i.e. seconds). In our FLOPs profiling analysis, we also include the models TerraMind [29] and Prithvi-EO-2.0 [23] (see Sec. V-E).
@@
Sentinel-2, L2A data. For the models we develop, implement, train, and evaluate that are based on the EO FM PhilEO [3], we use images of size 128128x128128, and also $10$ spectral bands (channels), i.e. multi-spectral satellite Sentinel-2 data. More specifically, as in [3], we use the $10$ and $20$ metre spatial resolution spectral bands from the satellite Sentinel-2999http://sentiwiki.copernicus.eu/web/s2-mission#S2Mission-SpatialResolutionS2-Mission-Spatial-Resolutiontrue.
@@ V-C Comparison of scaling-up for U-Net and ViT models
Jointly increasing the pretraining dataset size and the model size. In this paper, we focus not only on scaling-up the volume of the pretraining data (i.e. to 23TB) [15], but also on scaling the number of parameters and the architecture, also studying the transition from CNN-based U-Net models to ViT backbones combined with the UPerNet decoder, as well as to Mamba for images. Here, the recent Mamba S6 SSM is a potential effective paradigm-shift solution [30], as it combines global dependencies modeling and linear computational scaling. The self-attention mechanism of the ViT is not sensitive to the token order, while on the contrary, the selectivity mechanism of Mamba (i.e. the selective scan) is sensitive to the token order. Notably, it needs a large hidden state dimension as it has an exponentially decaying recall accuracy based on the distance from the current token [36], [68]. In addition, ViT has quadratic complexity with respect to the image token sequence length, due to the attention Query, Key, Value mechanism, which is based on the scaled dot product similarity measure, while the 2D Mamba S6 SSM model is faster and more efficient as it achieves linear complexity with respect to the image token sequence length [30], [31].
@@
Figure 18: Road density pixel-wise regression in RMSE for the models ViT and Mamba. Comparison to the TerraMind [29] and Prithvi-EO-2.0 [23] models. In the plot, the green line (Prithvi-EO-2.0) is below the pink line (TerraMind).
@@ V-D Comparison of scaling-up for Mamba and ViT models
Finally, because Mamba has benefits in efficiency, in addition to Figs. 8 and 9, we further examine the number of FLOPs of the Mamba SSM models. The Mamba UPerNet P16 model (i.e. from the preceding paragraph) has 100100M parameters and 12.6812.68G FLOPs. RS3Mamba has 168168M parameters and 18.6418.64G FLOPs. These two are significantly lower FLOPs than ViT UPerNet 100100M (i.e. see Sec. V-E, as well as Figs. 15-17) which has 388.51388.51G FLOPs, i.e. approximately $30$ times lower and this is because of the patch size $16$ (instead of $4$, i.e. larger patch sizes) and also because of Mamba selective scan linear complexity rather than ViT self-attention quadratic complexity.
@@ VI Conclusion
In this paper, we also compared with Mamba models (see Sec. III-C, as well as Sec. V-D), and this also one of our main contributions. Furthermore, we think our FLOPs profiling analysis, comparison, and benchmarking (i.e. in Sec. V, as well as in Figs. 8 and 9) opens the road to fairly evaluating different models (including GFMs) and will inspire others to adopt and use this real-world setting. Finally, it was a major achievement to process the 23TB of MajorTom on 32 cores of the Leonardo supercomputer davinci-1, in many days. This would have taken many more days, and even more months, on a single GPU. As future work, we plan to use JEPA (see Sec. IV), as well as to add multi-temporal capability (i.e. multiple time steps) [58], and multi-modality. Regarding the latter, we will add S-1 in addition to S-2 in the pretraining and not only in the S12-PhilEOBench (i.e. see Sec. III-D). As future work, we will also use Mamba 2D models with larger images, for example 512512x512512 (or even 10241024x10241024), as well as perform knowledge distillation of models to the same number of FLOPs (e.g., for On-Board AI) and compare and benchmark. Finally, as future work, we plan to also examine, compare with, and try to include in the model pipeline the recent AlphaEarth Foundations Satellite Embeddings [^10] http://deepmind.google/discover/blog/alphaearth-foundations-helps-map-our-planet-in-unprecedented-detail/. These are geo-located embeddings at 1010m resolution, and they also need the longitude and latitude information to be used in the downstream tasks.